What is the semi-annual coupon bond’s nominal yield to maturity (YTM), if the years to maturity is 15 years, and sells for 119% with coupons rate of 10%? Assume the par value of the bond is $1,000.


1
Expert's answer
2020-11-05T06:34:06-0500

"P=1.19\times1000=1190"

"1190=\frac{50}{(1+r)^1}+\frac{50}{(1+r)^2}+\frac{50}{(1+r)^3}+\frac{50}{(1+r)^4}+.....+\frac{1050}{(1+r)^{30}}"

r=0.0391283

"YTM=(1+ 0.0391283)^2-1)\times100=7.978"